Find the extreme values of z

 on the surface 2x

2

+3y

2

+z

2

−12xy+4xz=35



1
Expert's answer
2021-12-07T12:10:33-0500

A general concept: if f(x,y,z(x,y))=0 then differentiating with respect to x and y gets:

"f_x+f_zz_x = 0"

"f_y+f_zz_y = 0"

when "z_x = z_y = 0" , you'll get "f_x = f_y = 0."

So solve for the partials of f with respect to x and y are 0.

"f_x = 4x-12y+4z = 0"

"f_y = 6y-12x = 0"

solving for x,y as functions of z:

"x=z\/5,y=2z\/5"


Plug those into the original:

"2z^2\/25+12z^2\/25+z^2-24z^2\/25+8z^2\/25=35"

"23z^2=875"

"z=\pm 6.17"